OpenStudy (tennistar):

Nakim simplified 3 times the square root of 2x plus x times the square root of 8x minus 5 times the square root of 18x and got -10x times the square root of 2x for an answer. Part 1: Using complete sentences, explain what Nakim did wrong. Part 2: Show all your work to simplify the expression. (You can use "sqrt()" to show a square root. For example, 7 times the square root of a can be written as 7sqrt(a). )

OpenStudy (tennistar):

Please help me?

OpenStudy (tennistar):

@satellite73

OpenStudy (tennistar):

@myininaya

OpenStudy (tennistar):

@ParthKohli

OpenStudy (ganpat):

\[3(\sqrt{2x}) + x \sqrt{(8x - 5)}\] is this the basic equation ?? m little confused..

OpenStudy (ganpat):

wait.. i missed something, now i get it

OpenStudy (tennistar):

no\[3\sqrt{2x}+x \sqrt{8x}-5\sqrt{18x}\]

OpenStudy (tennistar):

\[ -10x \sqrt{2x}\] He got this

OpenStudy (ganpat):

\[3\sqrt{2x} + x \sqrt{8x} - 5\sqrt{18x} = -10x \sqrt{2x}\]

OpenStudy (tennistar):

is that correct?

OpenStudy (ganpat):

follow my steps.. i will explain you, consider left equation.. 1st term.. is fine

OpenStudy (tennistar):

as is

OpenStudy (ganpat):

now, consider 2nd term.. simplify it \[x \sqrt{(2 * 2 * 2 ) x}\] take 4 outside.. so sq rt of 4 = 2 so, what we get is \[2x \sqrt{2x}\]

OpenStudy (ganpat):

now take the third term, and solve similarly.. \[5 \sqrt{18x} = 5\sqrt{(3 * 3* 2) x} = 5 * 3 \sqrt{2x} = 15 \sqrt{2x}\]

OpenStudy (ganpat):

now put them back together and solve.. \[3\sqrt{2x} + 2x \sqrt{2x} - 15\sqrt{2x} = (3 -15) \sqrt{2x} + 2x \sqrt{2x} = -12\sqrt{2x} + 2x \sqrt{2x}\].. is the answer his mistake was he added the above two terms..

OpenStudy (ganpat):

sorry, i went on too lengthy.. hope u understood this..

OpenStudy (tennistar):

No, thank you so very much!!

OpenStudy (ganpat):

no mention.. just make sure u understand that.... 2 different terms cannot be added together.. like 2 and 2x .. as we saw above...